OpenStudy (anonymous):

Give equation in quadratic form, solve: 2^2x +21=10(2^x)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2^{2x} +21=10(2^x) }\) \(\normalsize\color{blue}{ 2^{2} \times 2^x +21=10(2^x) }\) \(\normalsize\color{blue}{ 4 \times 2^x +21=10(2^x) }\) \(\normalsize\color{blue}{ 4(2^x) +21=10(2^x) }\) \(\normalsize\color{blue}{ 21=10(2^x)-4(2^x) }\) \(\normalsize\color{blue}{ 21=6(2^x) }\) \(\normalsize\color{blue}{ 21/6=2^x }\)

OpenStudy (solomonzelman):

take log of both sides and solve for x.